please solve
4 1/3=(1/2)t

Answer by TutorDelphia(193)

Lets convert 4 1/3 to an improper fraction to start. Multiply 4*3 and add 1
13/3=(1/2)t
We can either multiply by 2 or divide by 1/2 (and to divide by 1/2 we flip and multiply so we end up multiplying by 2/1 which is 2 anyway)
26/3=t
Now just divide 26 by 3 and we ge 8 with a remainder of 2 so the answer is
8 2/3=t